STP Corp. uses no debt. The weighted average cost of capital is 10.4 percent. The current market value of the equity is $18 million and the corporate tax rate is 35 percent.

  

What is the EBIT? (Enter your answer in dollars, not millions of dollars, i.e. 1,234,567. Round your answer to 2 decimal places. (e.g., 32.16))

  

  EBIT $   
Answer #1

Current market value of equity=(EBIT(1-tax rate))/WACC

18million=(EBIT(1-0.35))/0.104

18,000,000*0.104=EBIT*0.65

EBIT=1872000/0.65

which is equal to

=$2,880,000.